The perimeter of a rectangle is 106 m. The length is 2m more than twice the width. find the dimension
Answer by oberobic(2304)

P = 106
P = 2L + 2W
.
L = 2W +2
2W +2 = L
.
substitute
.
2(2W+2) + 2W = 106
4W +4 +2W = 106
6W = 102
W = 17
.
L = 2W+2
L = 2*17 + 2
L = 36
.
Answer: Length = 36, Width = 17
.
Check the perimeter
2*36 =72
2*17 =34
72+34 = 106
Correct.
.
Done.